OpenStudy (anonymous):

1. The problem statement, all variables and given/known data From Elementary Differential Equation by Boyce and Diprima Chapter 2 Miscellaneous Problems #11 (x^2+y)dx + (x+e^x)dy = 0 ANS:(x^3/3)+xy+e^x=c 2. Relevant equations multiplying an integrating factor to make the DE exact: 1. du/dx = u(My - Nx)/ N 2. du/dx = u(Nx-My)/ M 3. The attempt at a solution First try: I guessed this can be changed into exact DE so, I tried with the two above equation: equation 1 gave me: du/u = e^x/(x+e^x) I don't know how to solve this... then equation 2 gave me: u = e^((e^x)

OpenStudy (anonymous):

u = e^((e^x)*ln(x^2+y)) I am not sure if multiply this integrating factor to the original DE will make it exact... Second try: I manipulated the given DE and changed it to a linear form: dy/dx = -(x^2+y)/(x+e^x) dy/dx + 1/(x+e^x) * y = (-x^2)/(x+e^x) and I found integrating factor to be: I = e^∫1/(x+e^x) dx which I am unable to solve...

OpenStudy (anonymous):

Third try: take integral of both side ∫(x^2+y)dx + (x+e^x)dy = ∫0 (x^3)/3 + yx + xy + ye^x = c //move constant c from left side to right side (x^3)/3 + 2xy + ye^x = c however it's not quite the same as the answer... Wolframalfa does not solve this one!!!